OpenStudy (anonymous):

What is 4x + 2y = -20 written in slope-intercept form? A. y = 2x - 20 B. y = -4x + 20 C. y = 20x - 4 D. y = -2x - 10

hero (hero):

D

hero (hero):

4x + 2y = -20 2y = -20 - 4x y = (-20 - 4x)/2 y = - 10 - 2x y = -2x - 10